Eight times the sum of a number and -2 is the same as nine times the numbers. Find the number.
Answer by checkley71(8403)

8(x-2)=9X
8X-16=9X
8X-9X=16
-X=16
X=-16 ANSWER.
PROOF:
8(-16-2)=9*-16
8*-18=-144
-144=-144